# Joint distribution of leftmost digits in positional notation and Schanuels’s conjecture

### 03 Mar 2026 v1

@@ 3 Schanuel’s conjecture and its consequences
$\zeta_{i}(z_{1},...,z_{m})=\infty$ on an (m1)(m-1)-dimensional $\mathbb{C}$-subspaces $S_{i}$ of $\mathbb{C}^{m}\subset(\mathbb{C}\cup\{\infty\})^{m}.$ The hypothesis in conjecture 2 implies that $S_{1},...,S_{m}$ are distinct. Therefore $\zeta_{1},...,\zeta_{m}$ and hence $\frac{1}{\ln b_{1}},...,\frac{1}{\ln b_{1}}$ are rationally independent. This proves that conjecture 1 implies conjecture 2. The proof of the reverse implication is straightforward.
@@ 4 Joint distribution for $n\geq 3$ bases
If some pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then $d_{B}$ is not surjective. If no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then conjecture 1, so Schanuel’s conjecture, implies that $d_{B}$ is surjective.
@@
The first assertion follows from Theorem 1. Assume that no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent and assume conjecture 1. Then proposition 1 implies conjecture 2 which implies that the entries of $\omega$ are rationally independent. Therefore Lemma 2 implies that the image of $h:=\psi_{\omega}\circ\ln:\mathbb{R}_{+}\mapsto\mathbb{T}^{n}$ is dense in $\mathbb{T}^{n}.$ The same argument used in the proof of Theorem 1 shows that $d_{B}$ is surjective.

### 10 Mar 2026 v2

@@ 3 Schanuel’s conjecture and its consequences
$\zeta_{i}(z_{1},...,z_{m})=\infty$ on an (m1)(m-1)-dimensional $\mathbb{C}$-subspaces $S_{i}$ of $\mathbb{C}^{m}\subset(\mathbb{C}\cup\{\infty\})^{m}.$ The hypothesis in conjecture 2 implies that $S_{1},...,S_{m}$ are distinct. Therefore $\zeta_{1},...,\zeta_{m}$ and hence $\frac{1}{\ln b_{1}},...,\frac{1}{\ln b_{1}}$ are rationally independent.
@@ 4 Joint distribution for $n\geq 3$ bases
If some pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then $d_{B}$ is not surjective. If no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then conjecture 2, and hence Schanuel’s conjecture, implies $d_{B}$ is surjective.
@@
The first assertion follows from Theorem 1. Assume that no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent and assume conjecture 2. Then the entries of $\omega$ are rationally independent. Therefore Lemma 2 implies that the image of $h:=\psi_{\omega}\circ\ln:\mathbb{R}_{+}\mapsto\mathbb{T}^{n}$ is dense in $\mathbb{T}^{n}.$ The same argument used in the proof of Theorem 1 shows that $d_{B}$ is surjective.